OpenStudy (anonymous):

A bar of 39 inches is cut into two pieces such that twice the longer piece is 3 inches more than three times the shorter. How many inches is the longer piece?

OpenStudy (anonymous):

let the longer piece be of length x then shorter piec is of length is 39-x by the problem 2x-3=3(39-x) hope u can solve for x

OpenStudy (anonymous):

Thank you very much, Yes I get it :)))

OpenStudy (anonymous):

2x-3=117 -3x or 2x+3x=117+3 5x=120 x=120/5 x=24 hence the longer piece is of length 24 inches and the shorter one is of length 39-24=15 inches

OpenStudy (anonymous):

welcome dear